Solve using elimination by addition
2x+3y=1
3x-y=7

Answer by rfer(16322)

9x-3y=21
2x+3y=1
11x=22
x=2
-----------------
3*2-y=7
-y=1
y=-1
-----------------
2*2+3*-1=1
3*2--1=7